Shift report 20010831 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Aug 31 2001 - 16:13:40 EDT

  • Next message: Flemming Videbaek: "change in ZDCs and trigger setup"

    ------------------------------------------------------------
    Date: Fri Aug 31 16:13:40 2001, Shift: 08:00-16:00
    Supervisor: Alexandru JIPA, with: -
    Log book#8 pages 47 - 50
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beams up to 13.30 (dump).
    2. Runs 4521, 4522, 4524 (data, A polarity). Run 4523 tests(Flemming Videbaek)
    Run 4525 - 0 magnetic field (except D1). Run 4526, 4527 Test with different 
    triggers (FV).
    MRS moved to 40 degrees. D1-D4 have now B polarity. D5 has A polarity
    (500 A).
    3. No problems with detectors. Dana Beavis modified some positions at the ZDCs controls.
    Flemminf Videbaek solved some problems related to fb3 mentioned in Ramiro Debbe report's.
    Problems with D1 and D2 magnets after the polarity change and 
    currents settings. MCR called. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Aug 31 2001 - 16:14:10 EDT